# Novel techniques of imaging interferometry analysis to study gas and plasma density for laser-plasma experiments

### v1

@@ Abstract
Laser-plasma based experiments are always more demanding about the plasma characteristics which need to be generated during the interaction. This is valid for laser-plasma acceleration as well as for inertial confinement fusion experiments. Most of these experiments are moving toward high repetition rate operation regimes, making even more demanding the requests on the plasma sources and the diagnostics to be implemented.
@@ 1 Motivation
In imaging interferometry, the probe and the reference beams, usually originated by splitting the same laser beam, are recombined on a specific plane to create an interference pattern, called interferogram. The phase shift is made evident by the fringe deflection, which highlights the area where the different refractive index causes a phase shift in the probe. This method permits to have a quantitative measurement of the plasma electron/gas density and its spatial distribution. Still, the analysis of the raw data is a non-trivial task, prone to many sources of error and dependent on manual inputs whose consequences we tried to mitigate in other works [1]. In future laser-plasma experiments the repetition rate of the laser shots will be dramatically increased, creating a severe constraint on the imaging diagnostics and their analysis. High repetition rate requires tools able to furnish analyzed data in real-time with high affordability. At the present days, most of the used routines, based on the Fast Fourier Transform and Abel or other inversion algorithms [2, 3], require an action from the operator and even minor changes (e.g. in the region of interest of the image) require the operator to adjust the analysis parameters, resulting in time loss and potentially causing errors. This is particularly important for the characterization of laser-produced plasma density, which is crucial for the tuning of the laser properties and the target to reach the optimal laser-plasma interaction in inertial confinement fusion, high energy density research and laser-driven plasma wakefields accelerators.
In this article, we will show the progress in developing an automated tool for a fast and operator-independent analysis of the interferograms [4]. This could enable a real time feedback to help the operation of laser-plasma at high repetition rate. In this framework, we are developing a machine learning (ML) routine, which has been trained to fast analyse the interferograms and retrieve the corresponding phase shift. We will introduce the architecture we decided to use, the cGAN structure, and the methods we used to train the ML network. Some remarkable applications of this architecture are the translation of aerial images into maps, or the colorization of pictures [5]. Similar approaches have been developed for scientific data analysis (e.g. [6, 7, 8]). The program we have developed so far is able to retrieve the phase shift of a synthetic interferogram. We will discuss the further steps to improve it for the analysis of raw interferograms acquired in real experimental conditions. This is the first step toward the implementation of an automated tool to retrieve information from an interferogram in a fast and robust way for high repetition rate experiments.
@@ 2 Methods: Network and Dataset
To train and implement those networks, we used the open-source library TensorFlow, a specific tool developed for machine learning and artificial intelligence. This library can be handled through Python open source programming language and the interface Keras [9]. Different communities have successfully applied this approach to a large variety of image-to-image translation tasks, not only for scientific purposes [5, 6, 7, 8]. The high adaptability, the large community working on it and the possibility to retrieve information online is a significant advantage of this architecture.
Figure 1: Schematic of the cGAN architecture and the training workflow. During the ”training”, the input images are sent to the generator which produces output images. The discriminator received, either, a couple of input and output images, or a couple of input and target images. During the training, the discriminator is optimized to recognise the outputs as ”false” and the target as ”true”, while the discriminator is optimized to produce images the most similar possible to the targets. The optimization parameters are given by the ”loss functions”.
@@ 3 Training and preliminary results
We trained the cGAN over more than $4400$ input-target data. At the end of the training, we tested the network on a set of interferograms not used for the training. In Fig.3 there is an example of one of those tests. In the top left of the figure, there are the input image, i.e. the synthetic interferogram. The phase shift from which the interferogram is generated is shown in the top centre of the figure. The output phase shift generated by the generator is in the top-right. A comparison between the lineouts of the two phase shift images is present in the bottom line of the figure. The retrieved phase shift (normalized to $2\pi$) is quite noisy, but its variation is low compared to the maximum value of the dephasing. The lineouts of the target (blue) and generated (red) are shown in the bottom of the figure. The generated phase shift well reproduce the shape and the maximum value of the target image. The output of the discriminator is shown in the down-left image.

### v2

@@ Abstract
Laser-plasma based experiments are always more demanding about the plasma features which need to be generated during the interaction. This is valid for laser-plasma acceleration as well as for inertial confinement fusion experiments. Most of these experiments are moving toward high repetition rate operation regimes, making even more demanding the requests on the plasma sources and the diagnostics to be implemented.
@@ 1 Motivation
In imaging interferometry, the probe and the reference beams, usually originated by splitting the same laser beam, are recombined on a specific plane to create an interference pattern, called interferogram. The phase shift is made evident by the fringe deflection, which highlights the area where the different refractive index causes a phase shift in the probe. This method permits to have a quantitative measurement of the plasma electron/gas density and its spatial distribution. Still, the analysis of the raw data is a non-trivial task, prone to many sources of error and dependent on manual inputs whose consequences we tried to mitigate in other works [1]. In future laser-plasma experiments the repetition rate of the laser shots will be dramatically increased, creating a severe constraint on the imaging diagnostics and their analysis. High repetition rate requires tools able to furnish analyzed data in real-time with high affordability. At the present days, most of the used routines, based on the Fast Fourier Transform and Abel or other inversion algorithms [2, 3], require an action from the operator and even minor changes (e.g. the region of interest of the image) require the operator to adjust the analysis parameters, resulting in time loss and potentially causing errors. This is particularly important for the characterization of laser-produced plasma density, and it is crucial for the tuning of the laser and target properties to reach the optimal laser-plasma interaction in inertial confinement fusion, high energy density research and laser-driven plasma wakefields accelerators.
In this article, we will show the progress in developing an automated tool for a fast and operator-independent analysis of the interferograms [4]. This could enable a real time feedback to help the operation of laser-plasma at high repetition rate. In this framework, we are developing a machine learning (ML) routine, which has been trained to fast analyse the interferograms and retrieve the corresponding phase shift. We will introduce the architecture we decided to use, the cGAN structure, and the methods we used to train the ML network. This architecture has been already used for the translation of aerial images into maps, or the colorization of pictures [5]. Similar approaches have been developed for scientific data analysis (e.g. [6, 7, 8]). The program we have developed so far is able to retrieve the phase shift of a synthetic interferogram. We will discuss the further steps to improve it for the analysis of raw interferograms, acquired in real experimental conditions. This is the first step toward the implementation of an automated tool to retrieve information from an interferogram in a fast and robust way for high repetition rate experiments.
@@ 2 Methods: Network and Dataset
To train and implement those networks, we used the open-source library TensorFlow, a specific tool developed for machine learning and artificial intelligence. This library can be handled through Python, an open source programming language, and the interface Keras [9]. Different communities have successfully applied this approach to a large variety of image-to-image translation tasks, not only for scientific purposes [5, 6, 7, 8]. The high adaptability, the large community working on it and the possibility to retrieve information online is a significant advantage of this architecture.
Figure 1: Schematic of the cGAN architecture and the training workflow. During the ”training”, the input images are sent to the generator which produces output images. The discriminator received, either, a couple of input and output images, or a couple of input and target images. During the training, the discriminator is optimized to recognise the outputs as ”false” and the targets as ”true”, while the discriminator is optimized to produce images the most similar possible to the targets. The optimization parameters are given by the ”loss functions”.
@@ 3 Training and preliminary results
We trained the cGAN over more than 440044004400 input-target data. At the end of the training, we tested the network on a set of interferograms not used for the training. In Fig.3 there is an example of one of those tests. In the top left of the figure, there are the input image, i.e. the synthetic interferogram. The phase shift from which the interferogram is generated is shown in the top centre of the figure. The output phase shift generated by the generator is in the top-right. A comparison between the lineouts of the two phase shift images is present in the bottom line of the figure. The retrieved phase shift (normalized to $2\pi$) is quite noisy, but its variation is low compared to the maximum value of the dephasing. The lineouts of the target (blue) and generated (red) are shown in the bottom of the figure. As can be seen, the discrepancy with the original data is not particularly pronounced in the central region of the horizontal lineout, while it is more evident at the edges where slope of the phase shift is larger and goes rapidly to zero. The vertical lineout shows a good agreement between the maximum value of the target data and the one obtained from the trained generator. The output image correctly goes to zero, still having a small discrepancy at the lower values. This suggest that the network is less accurate when steeper slopes of density are analysed. Still, the generated phase shift well reproduce the shape and the maximum value of the target image and the effects of those discrepancies on the density measurement will be analysed in the next steps of this work.